OpenStudy (anonymous):

x - 2y = 6 3x - 6y = 18

OpenStudy (anonymous):

2x + 11 -2x + 11 2x - 11 -2x - 11

OpenStudy (texaschic101):

x - 2y = 6 ---> x = 2y + 6 now sub 2y + 6 in for x in the other equation and solve for y.

OpenStudy (anonymous):

So whats the answer foo

OpenStudy (anonymous):

i got 3 mins till test is over

OpenStudy (texaschic101):

3x - 6y = 18 3(2y + 6) - 6y = 18 6y + 18 - 6y = 18 6y - 6y = 18 - 18 0 = 0 infinite solutions

OpenStudy (texaschic101):

your welcome foo

OpenStudy (anonymous):

i think you're wrong because thats not in the choice

OpenStudy (texaschic101):

what are your choices

OpenStudy (anonymous):

2x + 11 -2x + 11 2x - 11 -2x - 11

OpenStudy (texaschic101):

oh...your looking for y ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

did you find it

OpenStudy (texaschic101):

x - 2y = 6 -->(-3) 3x - 6y = 18 ----------- -3x + 6y = -18 (result of multiplying by -3) 3x - 6y = 18 -------------add 0 - 0 = 0 there are infinite solutions let me try something else... 3x - 6y = 18 -6y = -3x + 18 y = -3x/-6 + 18/-6 y = 1/2x -3 I can't get any of your answer choices....

OpenStudy (anonymous):

Man you suck I'm done

hero (hero):

If we divide the second equation by 3 we get x - 2y = 6. This means both equations are the same.

hero (hero):

@nijalww, You posted the answer choices, but you didn't post the actual instructions of what to do. Don't get mad @texaschic101 , she's just trying to help.

OpenStudy (anonymous):

You suck too hero lolol and im kidding

OpenStudy (texaschic101):

are the answer choices wrong ?

hero (hero):

Possibly

OpenStudy (texaschic101):

I feel better.....that boy is rude